# A Time Series Model for Three Asset Classes used in Financial Simulator

### v1

@@ Abstract.
We create a dynamic stochastic general equilibrium model for annual returns of three asset classes: the USA Standard & Poor (S&P) stock index, the international stock index, and the USA Bank of America investment-grade corporate bond index. Using this, we made an online financial app simulating wealth process. This includes options for regular withdrawals and contributions. Four factors are: S&P volatility and earnings, corporate BAA rate, and long-short Treasury bond spread. Our valuation measure is an improvement of Shiller’s cyclically adjusted price-earnings ratio. We use classic linear regression models, and make residuals white noise by dividing by annual volatility. We use multivariate kernel density estimation for residuals. We state and prove long-term stability results.
@@ 1.1. Index levels and earnings
A classic way to find whether the stock market is priced fairly is to compare index levels with earnings. The classic cyclically adjusted price-earnings (CAPE) ratio, introduced by Robert Shiller, measures how reasonably the stock market (measured by the Standard & Poor 500 and its predecessors) is priced relative to fundamentals (earnings). It is computed by dividing the current index level by annual earnings, averaged over the previous 10 years. Thie latter, called cyclical averaging, is done to smooth out wild short-term fluctuations in earnings, and take the data from both recessions and expansions. Historically, this ratio is between 5 and 45. Its long-term median value is 17. If it is much larger than that, then the market is overpriced and is likely to crash soon, or at least not grow as fast as before. See the classic article [6] or the monograph [16], or a more recent article [12] and a textbook [10, Chapter 8, Sections 4, 6].
@@ 1.2. The new valuation measure
In our previous article [13], we modified the CAPE. We think that CAPE lost its importance because dividend payout decreased. Historically, about 70% of earnings of stocks were paid as dividends to shareholders. However, in the last few decades, this decreased to 40% on average. Instead of paying earnings as dividends, executives prefer to use earnings for stock buybacks. This increases the stock value.
Reasons are still disputed in the academic community, but likely include tax issues, [10, Chapter 8, Section 4], [1, Chapter 12, Section 5], as well as incentives for executives who get paid in bonuses of stocks in their own company, [1, Chapter 15]. During stock buybacks, overall earnings remain the same, but index levels increases. This, in turn, boosts the CAPE ration, but does not make the market fundamentally overpriced.
To capture this, in [13] we compared annual total returns with annual earnings growth. In our research, annual returns are geometric (not arithmetic, which are commonly quoted); total (including dividends); and nominal (not adjusted for inflation). Annual earnings growth is measured as logarithmic change, and also nominal (without inflation adjustments). The long-term difference is 4–5% per year. Cumulative total returns minus earnings have linear trend with slope 4–5%. After detrending, this becomes mean-reverting: Modeled as a classic autoregression of order 1. This detrended quantity is our new valuation measure. Our previous research showed significant predictive power for next year’s total returns.
@@
Figure 1. On the left: Shiller cyclically-adjusted price-earnings (CAPE) ratio on the log scale versus the new valuation measure proposed in [13]. They track each other before 2000, but after that they diverge. On the right: annual realized volatility for the Standard & Poor. This graph is taken from [13].
@@ 1.3. Market volatility
Assume we model earnings growth separately. Then we can model earnings and the index together, using this valuation measure. Indeed, this measure is the difference between total returns and earnings growth. Knowledge of earnings growth gives us knowledge of total returns. But it is much harder to model earnings growth terms: These are not independent identically distributed (IID), that is, not a (strong) white noise, judging by the autocorrelation function (ACF). In fact, the returns $Q(t)$ themselves have ACF close to zero, but the squared returns $Q^{2}(t)$ or their absolute values $|Q(t)|$ do not have zero ACF. For background, see the classic references [3, Chapter 7], especially Figures 7.1–7.3. A solution comes from our previous article [14], using the volatility of the S&P index. For background on such stochastic volatility models, see [3, Chapter 7, Section 4] or [9, Chapter 3, Section 1] (especially Figures 3.1, 3.2).
We used monthly average volatility index (VIX), computed daily by the Chicago Board of Options Exchange (CBOE) from implied volatility (computed using prices of European options traded on this exchange). See background in [1, Chapter 7, Section 2, Figure 7.3]; also [1, Chapter 10, Section 1] for volatility-driven strategies; or [10, Chapter 8, Section 4, Figure 8.5]. Dividing monthly S&P 500 returns by this VIX made these returns IID Gaussian. Unfortunately, applying this methodology directly to our setting is problematic: We have annual (not monthly) data, and the VIX data goes back only as far as 1990 (or 1986 for the version of VIX for S&P 100, a close relative of S&P 500).
@@ 1.4. Bond rates and spreads
However, there exist other factors influencing the S&P 500 total returns. Bond rates are very important. Bond and stock markets compete for the capital of investors. An increase in bond rates makes the bond market more attractive, thus reducing demand for stocks. As a benchmark, we took the BAA rate for US corporate bonds. This Moody’s rating is less than the perfect AAA, and is on the lower range for investment-grade bonds. Still, these bonds are relatively safe from defaults. We take average daily rate in December of 1927–2024. See the classic article [2] for reference. But here we take the BAA rate instead of the short-term Treasury rate, since it is easier to model it as an autoregression. Also, see textbooks [1, Chapter 9, Section 4] and [10, Chapter 9, Section 4].
@@
The Federal Reserve System has direct influence over short-term rates via its standard open-market operations (however, recently it sometimes intervenes in long-term bond market as well). When a recession strikes, the Federal Reserve lowers short-term rates. However, the long-term rate now combines the current short-term rate and the anticipated future short-term rates. Indeed, an alternative to buying long-term bonds is reinvesting many times in short-term bonds. In sum, when long-term rates are lower than the short-term rates, then a recession and a stock market crash are likely to happen soon. For the term spread, we consider December average daily data for 10 years minus 3 month Treasury rates. For reference, see for example [8], and more recent textbooks [1, Chapter 9, Section 3], especially Figure 9.9; or [10, Chapter 9].
@@ 1.5. Our main contributions
Long-term stability is a desirable property for any financial econometric model. We create a time series model which is called DSGE: dynamic (time series equations), stochastic (each time step includes random white noise, which is IID), and general equilibrium (stationary). See, for example, a textbook [7].
@@
The Excel data and Pyhon code for the fitting of the model, analysis of residuals, and the web app itself, is available in the repository asarantsev/sim Web app has back-end written in Python, using the Flask framework, with bare HTML for the front-end. The web address of the simulator itself is asarantsev.pythonanywhere.com
@@ 2.1. Preliminary remarks
For this background, see [9, Chapter 1, Section 1]. We consider total returns (including dividends or coupon payments, not just price changes), nominal (not inflation-adjusted) rather than real (inflation-adjusted), and geometric (log returns) rather than arithmetic: If 1\$ turned into 2\$ the returns are $\ln 2$ not 100%. Arithmetic and geometric returns $r_{a},r_{g}$ are related as $r_{g}=\ln(1+r_{a})$. However, if $r_{a},r_{g}$ are small, then $r_{a}\approx r_{g}$.
@@ 3. White Noise Analysis Methodology
For background on white noise and normality tests, see [3, Chapter 1, Section 6]. Also, for white noise testing, see [9, Chapter 1, Sections 2, 4], and for normality testing, see [9, Chapter 1, Section 5]; finally, see [9, Chapter 2, Section 6], and the result (2.10) there.
@@
Note that we perform the ACF computation both for $x_{1},\ldots,x_{N}$ and for $|x_{1}|,\ldots,|x_{N}|$. The necessity of doing this for $x_{1}^{2},\ldots,x_{N}^{2}$ and not just for $x_{1},\ldots,x_{N}$ is shown in [9, Chapter 3, Section 1] (especially Figures 3.1, 3.2), mentioned in the Introduction. To avoid computing higher moments, which might lead to increased standard error, we use absolute values instead of squares: $|x_{1}|,\ldots,|x_{N}|$. But both $x_{1},\ldots,x_{N}$ (the original values) and $|x_{1}|,\ldots,|x_{N}|$ (absolute values) give us the same critical values of the $L_{1}$ statistic.
We recall the standard time series results: Under the white noise hypothesis, we have the following weak convergence: $(\hat{\rho}(1),\ldots,\hat{\rho}(k))\Rightarrow\mathcal{Q}_{k}$, where $N\to\infty$, where $\mathcal{Q}_{k}$ is the standard $k$-variate normal distribution. From here, it is easy to derive asymptotic results for $L_{1}$ and $L_{2}$. But we decided to make Monte Carlo simulations and compute critical values by hand, since we are not sure about the rate of convergence. After all, this was in part what motivated the replacement of the standard $L_{2}$ Box-Pierce test with the Ljung-Box test, see the classic article [11].
@@ 4.1. Stock returns
It is reasonable to model domestic stock returns $Q$ by terms which are IID but not Gaussian. But after normalizing (dividing by volatility), $Q/V$ can be modeled as IID Gaussian, see Figure 3. See also Table 2.
Somewhat suprpisingly, we see it is reasonable to model international returns $I$ as IID Gaussian even without normalizing. But we think it is a statistical illusion because we do not have that much annual data for international stocks. We still decide to normalize international returns and consider $I/V$.
@@ 4.2. Stock market volatility
Following [14] with monthly implied volatility analysis, annual volatility on the log scale is modeled as an autoregression of order 1:
@@ 4.3. Bond rates and returns
Results are: $a_{B}=0.016611$ and $d_{B}=0.055884$. The parameter $m_{0}$ stands for the duration: sensitivity of corporate bond prices to changes in interest rates. Here, the duration is about $5.5$ years. The parameter $a_{B}$ measures default rates. The $p$-values for the Student $T$-test are less than $0.1\%$.
Finally, the BAA corporate bond rate is also modeled as an autoregression of order 1, after taking the logarithm to ensure they stay positive:
@@ 4.5. Including duration for stocks
$\Sigma_{5}:=\begin{bmatrix}1&-0.3&-0.12&-0.1&0.07\\ 0.3&1&-0.14&-0.04&-0.15\\ -0.12&-0.14&1&0.29&0.41\\ -0.1&-0.04&0.29&1&-0.07\\ 0.07&-0.15&0.41&-0.07&1\\ \end{bmatrix}$
@@ 5.1. The valuation measure
In this subsection, we mimic the econometric analysis of [13, Section 2.1]. Define the wealth process as $W(t)=\exp(Q(1)+\ldots+Q(t))$, with initial wealth $W(0)=1$. We use $L$-year averaged trailing earnings $\overline{E}$ for $L=10$, and their growth:
@@
Results of fit of (15) are shown in Table 4. Solving for $h,b,c$ in (13), (14), similarly to [13, (6)], we get:
@@ 5.3. Combined model of S&P returns and earnings
$\Sigma_{3}=\begin{bmatrix}1&-0.18&-0.42\\ -0.18&1&0.1\\ -0.42&0.1&1\\ \end{bmatrix}$
@@ 6.6. Proof of the main result
This is a well-known result, see for example the classic book [3]: Extending $\delta(t)$ for negative $t$, we solve
@@ 7.5. Retirement rules
In our previous article [13], we considered validity of the classic 4% withdrawal rate: A retired person withdraws 4% of the initial capital the first year, and then keeps withdrawal annual amount constant after adjusting for inflation. We found this rule works well. However, there we could not simulate the earnings growth directly. As discussed in the Introduction, it is an essential part of such wealth simulation. Thus, in [13] we needed to make certain assumptions on earnings growth, which we found reasonable.
In this article, we have a complete and closed model. We can simply simulate this 4% rule straightaway. As discussed above, we need to compensate for inflation by setting annual increases of withdrawals. We set 4% annual increase in withdrawals. We assume the constant stock/bond portfolio split, and the 50/50 split within stock portfolio between domestic and international stocks. We start with the current (as of June 2025) market conditions, so use the main (not advanced) version of the simulator.
@@ 8. Conclusions and Further Research
However, our research is quite limited. There exist many other more narrow asset classes. For the stocks, there are value/growth factors, small/middle/large size factors, sector/industry portfolios (IT, medical, utilities, railroads), countries (Canada, Germany, China). We can also combine these factors: For example, German small value stocks. For background, see textbooks [9, Chapter 6] or [1, Chapter 7, Section 3].
For the bonds, there are Treasury/corporate/municipal bonds, bonds of various maturities, corporate bonds with given ratings (say CCC), bonds from other countries, and their combinations: For example, junk foreign government bonds of other countries. See [10, Chapter 10, Section 2] or [1, Chapter 9].
We tried to extend our research for some of these asset classes, but failed:
  • Not enough data, sometimes going back only to 1990s and 2000s;
  • Regression residuals are not IID, judging by the ACF.

### v2

# Valuation Measure of the Stock Market using Stochastic Volatility and Stock Earnings

@@ Abstract.
We create a time series model for annual returns of three asset classes: the USA Standard & Poor (S&P) stock index, the international stock index, and the USA Bank of America investment-grade corporate bond index. Using this, we made an online financial app simulating wealth process. This includes options for regular withdrawals and contributions. Four factors are: S&P volatility and earnings, corporate BAA rate, and long-short Treasury bond spread. Our valuation measure is an improvement of Shiller’s cyclically adjusted price-earnings ratio. We use classic linear regression models, and make residuals white noise by dividing by annual volatility. We use multivariate kernel density estimation for residuals. We state and prove long-term stability results.
2020 Mathematics Subject Classification:
62P05, 91G15, 91G70. Keywords: autoregression, stationarity, white noise, stochastic volatility, price-earnings ratio, bond duration, multiple linear regression. Corresponding author: A.S. asarantsev@unr.edu ORCID: 0000-0002-7415-4892
@@ 1.1. Index levels and earnings
A classic way to find whether the stock market is priced fairly is to compare index levels with earnings. The classic cyclically adjusted price-earnings (CAPE) ratio, introduced by Robert Shiller, measures how reasonably the stock market (measured by the Standard & Poor 500 and its predecessors) is priced relative to fundamentals (earnings). It is computed by dividing the current index level by annual earnings, averaged over the previous 10 years. Thie latter, called cyclical averaging, is done to smooth out wild short-term fluctuations in earnings, and take the data from both recessions and expansions. Historically, this ratio is between 5 and 45. Its long-term median value is 17. If it is much larger than that, then the market is overpriced and is likely to crash soon, or at least not grow as fast as before. See the classic article [6] or the monograph [14], or a more recent article [11] and a textbook [9, Chapter 8, Sections 4, 6].
@@ 1.2. The new valuation measure
In our previous article [12], we modified the CAPE. We think that CAPE lost its importance because dividend payout decreased. Historically, about 70% of earnings of stocks were paid as dividends to shareholders. However, in the last few decades, this decreased to 40% on average. Instead of paying earnings as dividends, executives use earnings for stock buybacks. This increases the stock value.
Reasons are still disputed in the academic community, but likely include tax issues, [9, Chapter 8, Section 4], [1, Chapter 12, Section 5], as well as incentives for executives who get paid in bonuses of stocks in their own company, [1, Chapter 15]. During stock buybacks, overall earnings remain the same, but index levels increases. This boosts the CAPE ration, but does not make the market overpriced.
To capture this, in [12] we compared annual total returns with annual earnings growth. In our research, annual returns are geometric (not arithmetic, which are commonly quoted); total (including dividends); and nominal (not adjusted for inflation). Annual earnings growth is measured as logarithmic change, and also nominal (without inflation adjustments). The long-term difference is 4–5% per year. Cumulative total returns minus earnings have linear trend with slope 4–5%. After detrending, this becomes mean-reverting: Modeled as a classic autoregression of order 1. This detrended quantity is our new valuation measure. Our research showed significant predictive power for next year’s total returns.
@@
Figure 1. On the left: Shiller cyclically-adjusted price-earnings (CAPE) ratio on the log scale versus the new valuation measure proposed in [12]. They track each other before 2000, but after that they diverge. On the right: annual realized volatility for the Standard & Poor.
@@ 1.3. Market volatility
Assume we model earnings growth separately. Then we can model earnings and the index together, using this valuation measure. Indeed, this measure is the difference between total returns and earnings growth. Knowledge of earnings growth gives us knowledge of total returns. But it is much harder to model earnings growth terms: These are not independent identically distributed (IID), that is, not a (strong) white noise, judging by the autocorrelation function (ACF). In fact, the returns $Q(t)$ themselves have ACF close to zero, but the squared returns $Q^{2}(t)$ or their absolute values $|Q(t)|$ do not have zero ACF. For background, see the classic references [3, Chapter 7], especially Figures 7.1–7.3. A solution comes from using the volatility of the S&P index. For background on such stochastic volatility models, see [3, Chapter 7, Section 4] or [8, Chapter 3, Section 1] (especially Figures 3.1, 3.2).
We used monthly average volatility index (VIX), computed daily by the Chicago Board of Options Exchange (CBOE) from implied volatility (computed using prices of European options traded on this exchange). See background in [1, Chapter 7, Section 2, Figure 7.3]; also [1, Chapter 10, Section 1] for volatility-driven strategies; or [9, Chapter 8, Section 4, Figure 8.5]. Dividing monthly S&P 500 returns by this VIX made these returns IID Gaussian. Unfortunately, applying this methodology directly to our setting is problematic: We have annual (not monthly) data, and the VIX data goes back only as far as 1990 (or 1986 for the version of VIX for S&P 100, a close relative of S&P 500).
@@ 1.4. Bond rates and spreads
Other factors influence the S&P 500 total returns. Bond and stock markets compete for the capital of investors. An increase in bond rates makes the bond market more attractive, thus reducing demand for stocks. As a benchmark, we took the BAA rate for US corporate bonds. This Moody’s rating is less than the perfect AAA, and is on the lower range for investment-grade bonds. Still, these bonds are relatively safe from defaults. We take average daily rate in December of 1927–2024. See the classic article [2] for reference. For background, see textbooks [1, Chapter 9, Section 4] and [9, Chapter 9, Section 4].
@@
The Federal Reserve System has direct influence over short-term rates via its standard open-market operations. When a recession strikes, the Federal Reserve lowers short-term rates. However, the long-term rate now combines the current short-term rate and the anticipated future short-term rates. Indeed, an alternative to buying long-term bonds is reinvesting many times in short-term bonds. In sum, when long-term rates are lower than the short-term rates, then a recession and a stock market crash are likely to happen soon. For the term spread, we consider December average daily data for 10 years minus 3 month Treasury rates. For reference, see for example [7], and more recent textbooks [1, Chapter 9, Section 3], especially Figure 9.9; or [9, Chapter 9].
@@ 1.5. Our main contributions
The Excel data and Pyhon code for the fitting of the model, analysis of residuals, and the web app itself, is available in the repository asarantsev/sim Web app has back-end written in Python, using the Flask framework, with HTML for the front-end. The web address is asarantsev.pythonanywhere.com
@@ 2.1. Preliminary remarks
For this background, see [8, Chapter 1, Section 1]. We consider total returns (including dividends or coupon payments, not just price changes), nominal (not inflation-adjusted) rather than real (inflation-adjusted), and geometric (log returns) rather than arithmetic: If 1\$ turned into 2\$ the returns are $\ln 2$ not 100%. Arithmetic and geometric returns $r_{a},r_{g}$ are related as $r_{g}=\ln(1+r_{a})$. However, if $r_{a},r_{g}$ are small, then $r_{a}\approx r_{g}$.
@@ 3. White Noise Analysis Methodology
For background on white noise and normality tests, see [3, Chapter 1, Section 6]. Also, for white noise testing, see [8, Chapter 1, Sections 2, 4], and for normality testing, see [8, Chapter 1, Section 5]; finally, see [8, Chapter 2, Section 6], and the result (2.10) there.
@@
Note that we perform the ACF computation both for $x_{1},\ldots,x_{N}$ and for $|x_{1}|,\ldots,|x_{N}|$. The necessity of doing this not just for $x_{1},\ldots,x_{N}$ is shown in [8, Chapter 3, Section 1] (especially Figures 3.1, 3.2), mentioned in the Introduction. But both $x_{1},\ldots,x_{N}$ (the original values) and $|x_{1}|,\ldots,|x_{N}|$ (absolute values) give us the same critical values of the $L_{1}$ statistic.
We recall the standard time series results: Under the white noise hypothesis, we have weak convergence: $(\hat{\rho}(1),\ldots,\hat{\rho}(k))\Rightarrow\mathcal{Q}_{k}$, where $N\to\infty$, where $\mathcal{Q}_{k}$ is the standard $k$-variate normal distribution. From here, it is easy to derive asymptotic results for $L_{1}$ and $L_{2}$. But we decided to make Monte Carlo simulations and compute critical values by hand, since we are not sure about the rate of convergence. After all, this was in part what motivated the replacement of the standard $L_{2}$ Box-Pierce test with the Ljung-Box test, see the classic article [10].
@@ 4.1. Stock returns
It is reasonable to model domestic stock returns $Q$ by terms which are IID but not Gaussian. But after normalizing (dividing by volatility), $Q/V$ can be modeled as IID Gaussian, see Figure 3. See also Table 2. Suprpisingly, we see it is reasonable to model international returns $I$ as IID Gaussian even without normalizing. But we think it is a statistical illusion because we do not have that much annual data for international stocks. We still decide to normalize international returns and consider $I/V$.
@@ 4.2. Stock market volatility
Following our previous manuscript with Jihyun Park with monthly implied volatility analysis, annual volatility on the log scale is modeled as an autoregression of order 1:
@@ 4.3. Bond rates and returns
Results are: $a_{B}=0.016611$ and $d_{B}=0.055884$. The parameter $m_{0}$ stands for the duration: sensitivity of corporate bond prices to changes in interest rates. Here, the duration is about $5.5$ years. The parameter $a_{B}$ measures default rates. The $p$-values for the Student $T$-test are less than $0.1\%$. Finally, the BAA corporate bond rate is also modeled as an autoregression of order 1, after taking the logarithm to ensure they stay positive:
@@ 4.5. Including duration for stocks
$\mathbf{\Sigma}_{5}:=\begin{bmatrix}1&-0.3&-0.12&-0.1&0.07\\ 0.3&1&-0.14&-0.04&-0.15\\ -0.12&-0.14&1&0.29&0.41\\ -0.1&-0.04&0.29&1&-0.07\\ 0.07&-0.15&0.41&-0.07&1\\ \end{bmatrix}$
@@ 5.1. The valuation measure
In this subsection, we mimic the econometric analysis of [12, Section 2.1]. Define the wealth process as $W(t)=\exp(Q(1)+\ldots+Q(t))$, with initial wealth $W(0)=1$. We use $L$-year averaged trailing earnings $\overline{E}$ for $L=10$, and their growth:
@@
Results of fit of (15) are shown in Table 4. Solving for $h,b,c$ in (13), (14), similarly to [12, (6)], we get:
@@ 5.3. Combined model of S&P returns and earnings
$\mathbf{\Sigma}_{3}=\begin{bmatrix}1&-0.18&-0.42\\ -0.18&1&0.1\\ -0.42&0.1&1\\ \end{bmatrix}$
@@ 6.6. Proof of the main result
This is a well-known result, see the classic book [3]: Extending $\delta(t)$ for negative $t$, we solve
@@ 7.5. Retirement rules
In our previous article [12], we considered validity of the classic 4% withdrawal rate: A retired person withdraws 4% of the initial capital the first year, and then keeps withdrawal annual amount constant after adjusting for inflation. We found this rule works well. However, there we could not simulate the earnings growth directly. As discussed in the Introduction, it is an essential part of such wealth simulation. Thus, in [12] we needed to make certain assumptions on earnings growth.
In this article, we have a closed model. We can simulate this 4% rule. As discussed above, we need to compensate for inflation by setting annual increases of withdrawals. We set 4% annual increase in withdrawals. We assume the constant stock/bond portfolio split, and the 50/50 split within stock portfolio between domestic and international stocks.
@@ 8. Conclusions and Further Research
However, our research is quite limited. There exist many other more narrow asset classes. For the stocks, there are value/growth factors, size factors, sector portfolios (IT, medical, utilities, railroads), countries (Canada, Germany, China). We can also combine these factors: For example, German small value stocks. For background, see [8, Chapter 6] or [1, Chapter 7, Section 3].
For the bonds, there are Treasury/corporate/municipal bonds, bonds of various maturities, corporate bonds with given ratings (say CCC), bonds from other countries, and their combinations. See [9, Chapter 10, Section 2] or [1, Chapter 9].
We tried to extend our research for these asset classes, but failed: Not enough data, sometimes only available since the 1990s and 2000s; Regression residuals are not IID.
@@
9. Statements and Declarations
The authors declare no conflicts of interest. All data and code are on GitHub/asarantsev and the data are taken from publicly available sources. This research is not supported by any external funding. A.S. is grateful to the Department of Mathematics and Statistics for the supportive atmosphere for undergraduate research.